# OmniGen2: Exploration to Advanced Multimodal Generation

### v1

@@ 1 Introduction
Unified image generation has attracted significant attention in recent years [80; 92; 73; 49]. For instance, OmniGen [80] employs a streamlined Transformer architecture capable of addressing diverse image generation tasks without the need for additional plugins or preprocessors. The recent breakthroughs of models like Gemini-2.0-flash [23] and GPT-4o [54] have further demonstrated the vast potential of this field, and heralded a new paradigm in multimodal intelligence, shifting the frontier from specialized models [3; 34] towards powerful, unified systems. Chameleon [72] and Emu3 [76] utilize discrete autoregressive methods across all modalities, while the Janus series [78; 12] introduces dual image encoders specifically designed for understanding and generation tasks. Transfusion integrates both autoregressive and diffusion processes within a single Transformer framework. However, the range of image generation tasks supported by these models remains limited.
@@
Beyond the model, we also explore to tackle the foundational data and evaluation challenges that impede progress in this domain. On the one hand, we started by collecting and reorganizing relevant open-source datasets. However, most of them are plagued by inherent quality limitations, especially in tasks such as image editing and in-context generation. This limitation has substantially contributed to the significant performance gap between open-source and commercial models. To address this issue, we developed a comprehensive pipeline for generating image editing and in-context data from video sources, and we plan to release these datasets to the community. Additionally, we constructed reflective data for image generation through an iterative generative process, with the aim of incorporating the reasoning and reflection capabilities of LLMs into multimodal generation.
On the other hand, our extensive evaluation of OmniGen2 reveals its competitive performance across various task domains, including text-to-image (T2I) generation, image editing, and in-context generation. Notably, for the in-context generation task, there is currently a lack of well-established public leaderboards to systematically assess and compare the key capabilities of different models. Existing resources, such as DreamBench [64], are insufficient to capture the complexity and requirements of real-world scenarios. To mitigate this limitation, we introduce the OmniContext benchmark, which includes eight task categories specifically designed to evaluate consistency with respect to individuals, objects, and scenes. Experimental results show that OmniGen2 achieves state-of-the-art consistency performance among open-source models.
@@ 3.1 In-Context Data
The in-context image generation task [79; 82; 35; 71] focuses on extracting a visual concept—such as a specific object, identiy or individual—from input images and accurately reproducing it within newly generated images. This task, also known as subject-driven generation [64], parallels in-context learning in large language models: the image generation model produces personalized outputs in real time based solely on the provided context, without the need for additional fine-tuning. While in-context image generation has been extensively explored due to its broad range of applications, the community still faces a notable shortage of high-quality datasets tailored to this task.
@@ 3.2.1 Inpaint Data
Most existing editing datasets are constructed using inpainting methods; however, they exhibit certain drawbacks: (1) The quality of images used is low: both due to low resolution and degradation after inpainting. (2) Editing instructions are inaccurate: previous work predefines editing instructions and uses inpainting models to generate images based on these instructions, but the inpainting models have poor instruction-following capabilities, causing a mismatch between editing instructions and original-inpainted image pairs.
@@
Figure 7: Create image edit pairs from video dataset. We first filter out frames belonging to different scenes to ensure contextual consistency, and then remove frames that exhibit significant changes in viewpoint.
@@ 3.3.2 Reflection Data
Inspired by previous advances in test-time scaling and self-reflection of large language models [24; 29; 42], we further explore the integration of reflection capabilities into multimodal generation models and demonstrate how test-time scaling can enhance the quality of image generation. In this section, we focus on describing the construction of the reflection data for subsequent model fine-tuning. The reflection data comprise an interleaved sequence of text and images, beginning with a user instruction, followed by the multimodal model’s generated image and step-by-step reflections on the previous generated outputs. Each reflection addresses two key aspects: 1) an analysis of the deficiencies or unmet requirements in relation to the original instruction, and 2) proposed solutions to address the previous image’s limitation.
@@ 5.4 In-context Generation
A distinguishing feature of OmniGen2 is its capcity to perform in-context generation, and we show the Qualitative visualizations in Figure 12. We introduce the OmniContext benchmark to provide a comprehensive evaluation of existing model’s performance in this domain. OmniContext comprises eight subtasks, with overall scores for each subtask presented in Table 6. As the inaugural model evaluated on this benchmark, OmniGen2 establishes a strong baseline, achieving an overall score of 7.18. These results show OmniGen2’s proficiency in disentangling the subject’s identity from its original background and re-rendering it accurately according to new textual instructions. OmniGen2 exhibits significant improvements over competing models across all task types, demonstrating superior prompt-following ability and subject consistency.
The detailed metrics for each subtask are presented in Tables  7,  8 and  9. Several noteworthy observations can be drawn from these results. OmniGen2 consistently outperforms all open-source baselines across every evaluation metric, irrespective of whether the input comprises a single image or multiple images. Among closed-source models, GPT-4o [54] achieves the highest scores in both the Overall and Prompt Following metrics, while Flux.1 Kontext [35] demonstrates superior performance in Subject Consistency. In contrast, Gemini-2.0-Flash [23] exhibits comparatively weaker results across these benchmarks.
@@ 6 Limitations
(1) Performance Disparity Between English and Chinese Prompts. As shown in the first row of Figure 14, prompts in English generally yield better results than those in Chinese. For instance, when using a Chinese prompt, the generated image exhibits a minor inconsistency between input image and edited image.
(2) Limited Generalization to Certain Instructions. The second row highlights OmniGen2’s difficulty in modifying human body shapes, likely due to the scarcity of real-world data capturing such variations.
(3) Sensitivity to Input Image Quality. As illustrated in Figure 14, the quality of the generated output is highly sensitive to the quality of the input image. When we input a low-quality image (generated by adding noise to the raw image), the resulting images exhibit significant degradation, with details becoming notably blurred. Furthermore, downsampling the input image to a maximum dimension of 256 pixels leads to further loss of clarity and detail, and the model’s ability to accurately follow generation instructions is substantially reduced.
(4) Ambiguity in Multi-Image Inputs. The third row of Figure 14 demonstrates that the model’s performance improves when the prompt explicitly specifies the correspondence between objects and their source images (e.g., “the bird from image 1 and the desk from image 2”), indicating a sensitivity to ambiguous multi-source instructions.
@@ 7.1 Image Generation
Diffusion models have achieved remarkable success in generative modeling, particularly for high-fidelity image synthesis [18; 9; 15]. Pioneering systems such as the Stable Diffusion (SD) series [62; 56; 18], DALL·E [60], and Imagen [28] have demonstrated strong text-to-image capabilities, laying the foundation for controllable and scalable generation. To enhance controllability, models like ControlNet [89] and T2I-Adapter [50] introduce external condition modules, while StyleShot [19], InstructPix2Pix [5], and EMU-Edit [67] focus on fine-grained editing and instruction-driven generation. These developments highlight the increasing flexibility of diffusion models in various conditional settings. Recently, a growing body of work has shifted attention toward unified image generation, which aims to handle various image generation tasks within a single model [80; 49; 11; 73]. OmniGen jointly models text and images within a single transformer to achieve unified image generation. UniReal treats various tasks as video generation prediction, and designs different prompts for task differentiation.
Recently, several studies have introduced Chain-of-Thought reasoning to enhance image generation tasks [25; 74; 30]. However, research on reflection mechanisms in this context remains limited. ReflectionFlow [95] proposes a reflection framework that integrates multiple components, including the original text-to-image (T2I) model, a fine-tuned multimodal large language model (MLLM) for instruction generation, and an image editing model. Notably, ReflectionFlow requires training only on single instruction-image pairs. In contrast, OmniGen2 explores an integrated approach, enabling reflection and subsequent image regeneration within a single model. Moreover, OmniGen is fine-tuned on multi-turn reflection data rather than single-turn editing data, allowing it to better capture the iterative nature of real-world image editing scenarios.

### v2

@@ 1 Introduction
Unified image generation has attracted significant attention in recent years [80; 92; 73; 48]. For instance, OmniGen [80] employs a streamlined Transformer architecture capable of addressing diverse image generation tasks without the need for additional plugins or preprocessors. The recent breakthroughs of models like Gemini-2.0-flash [22] and GPT-4o [53] have further demonstrated the vast potential of this field, and heralded a new paradigm in multimodal intelligence, shifting the frontier from specialized models [3; 33] towards powerful, unified systems. Chameleon [72] and Emu3 [76] utilize discrete autoregressive methods across all modalities, while the Janus series [78; 11] introduces dual image encoders specifically designed for understanding and generation tasks. Transfusion integrates both autoregressive and diffusion processes within a single Transformer framework. However, the range of image generation tasks supported by these models remains limited.
@@
Beyond the model, we also explore to tackle the foundational data and evaluation challenges that impede progress in this domain. We started by collecting and reorganizing relevant open-source datasets. However, most of them are plagued by inherent quality limitations, especially in tasks such as image editing and in-context generation. This limitation has substantially contributed to the significant performance gap between open-source and commercial models. To address this issue, we developed a comprehensive pipeline for generating image editing and in-context data from video sources, and we plan to release these datasets to the community. Additionally, we constructed reflective data for image generation through an iterative generative process, with the aim of incorporating the reasoning and reflection capabilities of LLMs into multimodal generation.
Our extensive evaluation of OmniGen2 reveals its competitive performance across various task domains, including text-to-image (T2I) generation, image editing, and in-context generation. Notably, for the in-context generation task, there is currently a lack of well-established public leaderboards to systematically assess and compare the key capabilities of different models. Existing resources, such as DreamBench [64], are insufficient to capture the complexity and requirements of real-world scenarios. To mitigate this limitation, we introduce the OmniContext benchmark, which includes eight task categories specifically designed to evaluate consistency with respect to individuals, objects, and scenes. Experimental results show that OmniGen2 achieves state-of-the-art consistency performance among open-source models.
@@ 3.1 In-Context Data
The in-context image generation task [79; 82; 34; 71] focuses on extracting a visual concept—such as a specific object, identiy or individual—from input images and accurately reproducing it within newly generated images. This task, also known as subject-driven generation [64], parallels in-context learning in large language models: the image generation model produces personalized outputs in real time based solely on the provided context, without the need for additional fine-tuning. While in-context image generation has been extensively explored due to its broad range of applications, the community still faces a notable shortage of high-quality datasets tailored to this task.
@@ 3.2.1 Inpaint Data
Although most existing editing datasets are constructed through inpainting techniques, they suffer from two primary limitations: (1) substandard image quality, stemming from both inherent low resolution and post-processing degradation during inpainting. (2) Editing instructions are inaccurate: previous work predefines editing instructions and uses inpainting models to generate images based on these instructions, but inpainting models have poor instruction-following capabilities, causing a mismatch between editing instructions and original-inpainted image pairs.
@@
Figure 7: Create image edit pairs from videos. We first filter out frames belonging to different scenes to ensure contextual consistency, and then remove frames that exhibit significant changes in viewpoint.
@@ 3.3.2 Reflection Data
Inspired by previous advances in test-time scaling and self-reflection of large language models [23; 28; 41], we further explore the integration of reflection capabilities into multimodal generation models and demonstrate how test-time scaling can enhance the quality of image generation. In this section, we focus on describing the construction of the reflection data for subsequent model fine-tuning. The reflection data comprise an interleaved sequence of text and images, beginning with a user instruction, followed by the multimodal model’s generated image and step-by-step reflections on the previous generated outputs. Each reflection addresses two key aspects: 1) an analysis of the deficiencies or unmet requirements in relation to the original instruction, and 2) proposed solutions to address the previous image’s limitation.
@@ 5.4 In-context Generation
A distinguishing feature of OmniGen2 is its capacity to perform in-context generation, and qualitative visualizations are shown in Figure 12. We introduce the OmniContext benchmark to provide a comprehensive evaluation of the performance of the existing model in this domain. OmniContext comprises eight subtasks, with overall scores for each subtask presented in Table 6. As the inaugural model evaluated on this benchmark, OmniGen2 establishes a strong baseline, achieving an overall score of 7.18. These results show OmniGen2’s proficiency in disentangling the subject’s identity from its original background and re-rendering it accurately according to new textual instructions. OmniGen2 exhibits significant improvements over competing models in all types of tasks, demonstrating superior prompt-following ability and subject consistency.
The detailed metrics for each subtask are presented in Tables  7,  8 and  9. Several noteworthy observations can be drawn from these results. OmniGen2 consistently outperforms all open-source baselines in all evaluation metrics, regardless of whether the input comprises a single image or multiple images. Among closed-source models, GPT-4o [53] achieves the highest scores in both the Overall and Prompt Following metrics, while Flux.1 Kontext [34] demonstrates superior performance in Subject Consistency. In contrast, Gemini-2.0-Flash [22] exhibits comparatively weaker results on these benchmarks.
@@ 6 Limitations
1) Performance Disparity Between English and Chinese Prompts. As shown in the first row of Figure 14, prompts in English generally yield better results than those in Chinese. For instance, when using a Chinese prompt, the generated image exhibits a minor inconsistency between input image and edited image.
2) Limited Generalization to Certain Instructions. The second row highlights OmniGen2’s difficulty in modifying human body shapes, likely due to the scarcity of real-world data capturing such variations.
3) Sensitivity to Input Image Quality. As illustrated in Figure 14, the quality of the generated output is highly sensitive to the quality of the input image. When we input a low-quality image (generated by adding noise to the raw image), the resulting images exhibit significant degradation, with details becoming notably blurred. Furthermore, downsampling the input image to a maximum dimension of 256 pixels leads to further loss of clarity and detail, and the model’s ability to accurately follow generation instructions is substantially reduced.
4) Ambiguity in Multi-Image Inputs. The third row of Figure 14 demonstrates that the model’s performance improves when the prompt explicitly specifies the correspondence between objects and their source images (e.g., “the bird from image 1 and the desk from image 2”), indicating a sensitivity to ambiguous multi-source instructions.
5) In in-context generation tasks, the model occasionally fails to perfectly reproduce objects from the provided context. Increasing the guidance scale of image can partially alleviate this issue; however, it does not offer a complete solution. We hypothesize that significant improvements on such complex tasks may require further scaling of the model size.
@@ 7.1 Image Generation
Diffusion models have achieved remarkable success in generative modeling, particularly for high-fidelity image synthesis [17; 8; 14]. Pioneering systems such as the Stable Diffusion (SD) series [62; 56; 17], DALL·E [60], and Imagen [27] have demonstrated strong text-to-image capabilities, laying the foundation for controllable and scalable generation. To enhance controllability, models like ControlNet [89] and T2I-Adapter [49] introduce external condition modules, while StyleShot [18], InstructPix2Pix [4], and EMU-Edit [67] focus on fine-grained editing and instruction-driven generation. These developments highlight the increasing flexibility of diffusion models in various conditional settings. Recently, a growing body of work has shifted attention toward unified image generation, which aims to handle various image generation tasks within a single model [80; 48; 10; 73]. OmniGen jointly models text and images within a single transformer to achieve unified image generation. UniReal treats various tasks as video generation prediction, and designs different prompts for task differentiation.
Recently, several studies have introduced Chain-of-Thought reasoning to enhance image generation tasks [24; 74; 29]. However, research on reflection mechanisms in this context remains limited. ReflectionFlow [95] proposes a reflection framework that integrates multiple components, including the original text-to-image (T2I) model, a fine-tuned multimodal large language model (MLLM) for instruction generation, and an image editing model. Notably, ReflectionFlow requires training only on single instruction-image pairs. In contrast, OmniGen2 explores an integrated approach, enabling reflection and subsequent image regeneration within a single model. Moreover, OmniGen is fine-tuned on multi-turn reflection data rather than single-turn editing data, allowing it to better capture the iterative nature of real-world image editing scenarios.